OpenStudy (anonymous):

Coral created a painting that was 4 1/3 ft long and 1 1/5 t wide. What was the area of Coral's painting? Express your answer in simplest form. A. 6 2/5 ft² B.5 8/15 ft² C.5 1/5 ft² D. 4 1/15 ft²

OpenStudy (jdoe0001):

 \(\bf \textit{area of a rectangle}=length\cdot width\to 4\frac{1}{3}\cdot 1\frac{1}{5}\)

OpenStudy (anonymous):

So which answer chocie is it

OpenStudy (jdoe0001):

well... what did you get for the multiplication?

OpenStudy (anonymous):

* choice